Probiem &.8-Enhanced- with MacBook Air 2 5 7 9
Answer #1

The image is not very clear. I had a difficulty in reading out the values....

from what I see, the angle is 21 degree , mass is 30 Kg , coefficient of friction is not clear. However, here is the method how to solve for required force.

Summing up the forces in x direction, we have

Fcos\theta - uN = 0

In y direction

N = mg + Fsin\theta

Put this N in first equation

Fcos\theta - u (mg + Fsin\theta) = 0

Therefore,

F = \frac{\mu mg}{cos\theta -\mu sin\theta }

Just put the values....here \mu is coefficient of friction, m is mass , g is 9.8, \theta = 21